# Task Me Anything

### v1

@@ F.1 Experiment details
F.2 Experiments on approximations under different budgets.
To evaluate the performance of approximation algorithms under different budgets, we conducted an experiment using Qwen-VL-Chat as the target model on 2D how-many tasks. We tested three query approximation algorithms on four types of queries: Top-K query, Threshold query, Model comparison query, and Model debugging query. The experiments were performed under budgets of $1,000$, $2,000$, and $3,000$. The results of the experiment can be found in Table 171819, and 20.
@@
F.3 Query results approximation experiments in ImageQA
@@
F.4 Query results approximation experiments in VideoQA
@@
Appendix G Details of Analysis and Case Study
G.1 What task metadata are models good or bad at?
@@
G.2 How do small models compare against large models? (continued)
@@
Further, upon a closer look with our interface, we identify a few relations where InstructBLIP-7B outperforms InstructBLIP-13B by a large margin e.g. 50% (Figure 35). Similarly, we also retrieve a few actions and objects where Video-LLaMA-2-7B performs much better e.g. by 20% than Video-LLaMA-2-13B (Figures 36 and 37).
@@
G.3 Do Task-Me-Anything yield results similar to existing benchmarks?
@@
Appendix H Datasheet for Task-Me-Anything-Random
H.1 Motivation
@@ H.6 Maintenance
Appendix I Task Generator Cards

### v2

@@ G.1 Experiment details
G.2 Experiments on approximations under different budgets.
To evaluate the performance of approximation algorithms under different budgets, we conducted an experiment using Qwen-VL-Chat as the target model on 2D how-many tasks. We tested three query approximation algorithms on four types of queries: Top-K query, Threshold query, Model comparison query, and Model debugging query. The experiments were performed under budgets of $1,000$, $2,000$, and $3,000$. The results of the experiment can be found in Table 252627, and 28.
@@
G.3 Query results approximation experiments in ImageQA
@@
G.4 Query results approximation experiments in VideoQA
@@
Appendix H Details of Analysis and Case Study
H.1 What task metadata are models good or bad at?
@@
H.2 How do small models compare against large models? (continued)
@@
Further, upon a closer look with our interface, we identify a few relations where InstructBLIP-7B outperforms InstructBLIP-13B by a large margin e.g. 50% (Figure 39). Similarly, we also retrieve a few actions and objects where Video-LLaMA-2-7B performs much better e.g. by 20% than Video-LLaMA-2-13B (Figures 40 and 41).
@@
H.3 Do Task-Me-Anything yield results similar to existing benchmarks?
@@
Appendix I Datasheet for Task-Me-Anything-Random
I.1 Motivation
@@ I.6 Maintenance
Appendix J Task Generator Cards